Find the standardized test statistic to test the claim that μ1 < μ2. Two samples are random, independent, and come from populations that are normally distributed. The sample statistics are given below. Assume that σ 2 /1 = σ 2 /2 .

n1 = 15 n2 = 13 x1 = 27.88 x2 = 30.43 s1 = 2.9 s2 = 2.8
Answer #1

Solution:

Since the population variances are unknown , we use t test.

Since , Assume that σ 2 /1 = σ 2 /2 case , so we use pooled variance .

The test statistic t is given by

t = T1 T2 + ) 111 112

where   s_{p}^{2} denotes the pooled variance.

Therefore ,

t = \frac{\bar x_{1}-\bar x_{2}}{\sqrt{ \frac{(n_{1}-1)s_{1}^2 + (n_{2}-1)s_{2}^2}{ n_{1}+n_{2}-2}(\frac{1}{ n_{1}}+\frac{1}{ n_{2}}) }}

= 27.88 - 30.43 (15-12.92 +(13-1/2.82 15-13—2 15 高) 13

= -2.35766

= -2.358

The standardized test statistic is

t = -2.358